A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 12.0 cm , giving it a charge of -10.0 μC .

Part A

Find the electric field just inside the paint layer.

Part B

Find the electric field just outside the paint layer.

Part C

Find the electric field 5.50 cm outside the surface of the paint layer.
Answer #1

Part A:

Electric field just inside = zero (as Gaussian surface will enclose no charge)

Part B:

The electric field just outside, we can assume the sphere as point charge,

E = kQ/r^2 = -9*10^9*10*10^-6/0.06^2 = -25000000 N/C

Part C:

Electric field at 11.5 cm from the center,

E= kQ/d^2 = -9*10^9*10*10^-6/0.115^2 = -6805293.01 N/C